OpenStudy (anonymous):

v^2=u^2+2as u=3 a=9.8 s=12 Work out the value of v. Give your answer correct to 1d.p.

OpenStudy (anonymous):

@hazem-wardat @bibby

OpenStudy (bibby):

1 DP means we're rounding to the tenths place. Just repeat what we did last question. solve for u^2 + 2as and take the square root of both sides

OpenStudy (anonymous):

so is this 9+2*9.8*12

OpenStudy (anonymous):

or 9+2*9.8+12

OpenStudy (bibby):

\[\large 9+ 2*a*s = 9 + (2*9.8*12)\]

OpenStudy (anonymous):

244.2?

OpenStudy (bibby):

it'll be clearer to you if you line up the equations and then fill in the variables. \[\huge u^2+2as\]\[\huge 3^2+[2*12*9.8]\]

OpenStudy (bibby):

yep!

OpenStudy (anonymous):

then do i square root 244.2

OpenStudy (bibby):

you do. because you have \[\huge v^2 =244.2 \]but you want v alone

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

sorry i took so long the doctor had to check my temp again

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

lol are you in a doctor's office? no worries

OpenStudy (anonymous):

No I have stayed in the hospital for a week now they say they will send me out next week

OpenStudy (anonymous):

When i am well hopefully

OpenStudy (bibby):

oh :<, feel better. Anyhow we take the square root of both sides v^2 = 244.2

OpenStudy (anonymous):

15.626899884494 is square root of 244.2

OpenStudy (anonymous):

15.6 to be short

OpenStudy (bibby):

Exactly. Re-read the directions. We want it to one decimal place.

OpenStudy (bibby):

yep good job

OpenStudy (anonymous):

thnx again

OpenStudy (anonymous):

but what is: s=ut=0.5at^2 a=9.8 t=5 u=7

OpenStudy (bibby):

ask it in another question

OpenStudy (anonymous):

okay